OpenStudy (anonymous):

You invest an initial $2,000 in an account that has an annual interest rate of 6%, compounded daily. How much money will you have in the account after 15 years? Round your answer to the nearest whole number.

OpenStudy (camerondoherty):

First, What is 0.06 times 15?

OpenStudy (camerondoherty):

Once you find the product of that, multiply it by 2,000 and add it to 2,000

OpenStudy (anonymous):

0.9

OpenStudy (anonymous):

32000

OpenStudy (anonymous):

\[y=yoe ^{tk}\]

OpenStudy (anonymous):

yo: initial amount 2000 t: time 15 k: rate 6/100 dealy

OpenStudy (kropot72):

\[\large A=2000(1+\frac{0.06}{365})^{(365\times15)}=2000(1.000164384)^{5475}\]